40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 17 20:22:07 2012, Jamie, Update, General, Vent plan: WE VENT TOMORROW 
    Reply  Mon Dec 17 21:33:45 2012, Manasa, Update, General, Vent plan: WE VENT TOMORROW 
    Reply  Tue Dec 18 07:03:40 2012, Manasa, Update, General, Vent plan: WE VENT TOMORROW 
    Reply  Tue Dec 18 10:43:20 2012, Steve, Update, VAC, vent has completed vent74.pngbvent.pngbefVent.png
       Reply  Wed Dec 19 11:53:23 2012, Steve, Update, VAC, AC removed workingatatm.png
       Reply  Thu Jan 24 14:40:09 2013, Steve, Update, VAC, RGA backgroung Screenshot.png
       Reply  Mon Feb 11 16:16:27 2013, Steve, Update, VAC, 55 days at atmoshere atm55d.png
Message ID: 7842     Entry time: Mon Dec 17 20:22:07 2012     Reply to this: 7843   7848   7849
Author: Jamie 
Type: Update 
Category: General 
Subject: Vent plan: WE VENT TOMORROW 

We will vent tomorrow (12/18)

After lengthy discussion, I have determined that we should vent now, with the primary goal of this vent to replace the input steering PZTs with the new active tip-tilts.  Since we still don't understand exactly what's going on with the PRC, it's unclear what we would do to attack the problem.  We need to do more modeling and measurements first.  We should limit the goal of this vent to replacing the PZTs, and then close up and do more measurements with better modeling and improved input point in hand.

There is limited time this week before everyone leaves for two-week holidays on Thursday or Friday.  The reason to not vent would be that we don't want to leave the IFO at air during the two week holiday.  People seem to think that this is not a problem, so we don't gain anything by waiting.  Therefore we vent now and do what we can before people take off.

The goal for this week is to replace PZT1 only:

Now: Jenne and Manasa are doing vent prep.  Manasa is lowering the input power and preparing the mode cleaner.  Jenne centered IPPOS and IPANG.  This will allow us to check how the input alignment changed during the vent.

12/18: Steve is going to start the vent as soon as he gets back from the dentist, at around 10am.  He will regulate the vent such that we are ready to lock the mode cleaner by 4pm.  At that time we will lock the MC and recheck the input alignment with IPPOS/ANG, with the idea being to see if anything moves during the vent.

12/19: First thing in the morning we take off the access connector ONLY.  The access connector is all we'll need to replace PZT1.    Put light door on the OMC chamber side immediately, since we won't need to get in there at all.  We won't need the light access connector.

For the rest of Wednesday we'll remove PZT1 and install the new active TT.  We'll be using the current out-of-vac cable for PZT1 for the new TT.  We should only have to modify the rack end of the cable to accommodate the coil driver.  This should be a small modification.  Given that we have no wiring diagrams we'll have to pin it out in situ.

12/20: Hopefully finish up TT installation.

Jenne leaves 12/20, Manasa and Jamie leave 12/21.  We will either leave light doors on access connector holes, or possibly Rana, Koji, and Steve will replace access connector on Friday so that we can pump down to 1 Torr or something so that we leave it there over the holiday.

After we return from vacation:

PZT2/TT2 installation.  This will be less straight forward since the new TT has a bigger foot print than PZT2 and will block the PRM optical levers.  We'll need one additional steering mirror to redirect the oplev around the TT.  See elog 7815.

Once the new TTs are installed, we'll reevaluate where we're at.  If PRC modeling has progressed and we have an idea of something to work on with the PRC, we can.  Otherwise, we'll just button up, pump down, and get on with some better PRC measurements.

ELOG V3.1.3-